five times the diffrence of a number and one yields the same result as twice the same number increase by three: five times the diffrence of a number and one yields the same result as twice the same number increase by three
Answer by ankor@dixie-net.com(4489)

Let x = "the number"
:
Write an equation for the statement:
"five times the difference of a number and one yields the same result as twice the same number increased by three"
5(x-1) = 2x + 3
:
5x - 5 = 2x + 3
:
5x - 2x = 3 + 5
:
3x = 8
x = 8/3
x = 22/3
;
:
Check solution using decimals
5(2.67 - 1) = 2(2.67) + 3
5(1.67) = 5.33 + 3
8.33 = 8.33